OpenStudy (anonymous):

t^2=6+8t solve the equation by completing the square and leave the irrational roots in simplest radical form. please help solve

OpenStudy (anonymous):

\[t^2-8t=6\] \[(t-4)^2=6+4^2=6+16=22\] \[t-4=\pm\sqrt{22}\] \[t=4\pm\sqrt{22}\]